Suppose that you invest in a two-year Treasury bond with a coupon rate of 6% and $1,000 par. Suppose that you buy this bond at a price of exactly $1,000. You intend to hold this bond to maturity and reinvest the coupons until the bond matures. You expect to reinvest the coupons in an account that pays an APR of 2.83%, with semi-annual compounding. What is the effective annual rate of return on your investment?
Answer #1

Since it is a Treasury bond, it is assumed that interest is paid semi annually.

Given,

Face Value of Bond (F) = $1,000. Term to maturity= 2 years.

Coupon Rate= 6% Therefore, Semi annual interest= $1,000*6%/2 = $30

The cash flows comprising the interest payments constitute an annuity at 2.83% interest (APR). Future value of the interest payments= $ 122.57 Calculated as follows:

F A B C D | | Ε G 1 Future Value of Annuity Payments at the end of each period 2 3 Future value of annuity is calculated usin

Total amount on redemption= Face Value + Future Value of interest stream

=$1,000 + $122.57 (FV)= $1,122.57

Purchase price (P)= $1,000

Effective annual rate of return=[(FV/P)^(1/n)]-1

Where FV= Total amount receivable on maturity, P= Purchase price and n= period of holding (number of years given as 2)

Substituting the values,

Effective annual rate of return= [(1,122.57/1,000)^(1/2)]-1 = 1.05951404 -1 = 5.951404%